The amount of trash generated by US households (in lbs per day) is Normally distributed. You a random sample of 8 people gives you an average of 10.00 lbs and a sample stdev of 5 lbs. What is the margin of error for a 99% Conf. Int. for the mean? (Leave your answer to two decimal places.)
Answer #1

Solution :

Given that,

t /2,df = 3.499

Margin of error = E = t/2,df * (s /n)

= 3.499 * (5 / 8)

Margin of error = E = 6.19